If one bag of candy costs 45cents and contains 25bars, how many tera tons of fat are in $128.00 worth of candy if each bar contains 96milligrams of fat?
Answer by checkley71(8403)

.45/12800=25*96/96X
.45/12800=25/X NOW CROSS MULTIPLY
.45X=320,000
X=320,000/.45
X=711,111.11 MGRAMS OF FAT FOR $128.00